Status Report

NASA Hubble Space Telescope Daily Report #4718

By SpaceRef Editor
October 18, 2008
Filed under , ,
NASA Hubble Space Telescope Daily Report #4718


Continuing to collect World Class Science

PERIOD COVERED: 5am October 16 – 5am October 17, 2008 (DOY 290/0900z-291/0900z)


FGS 11212

Filling the Period Gap for Massive Binaries

The current census of binaries among the massive O-type stars is seriously incomplete for systems in the period range from years to millennia because the radial velocity variations are too small and the angular separations too close for easy detection. Here we propose to discover binaries in this observational gap through a Faint Guidance Sensor SNAP survey of relatively bright targets listed in the Galactic O Star Catalog. Our primary goal is to determine the binary frequency among those in the cluster/association, field, and runaway groups. The results will help us assess the role of binaries in massive star formation and in the processes that lead to the ejection of massive stars from their natal clusters. The program will also lead to the identification of new, close binaries that will be targets of long term spectroscopic and high angular resolution observations to determine their masses and distances. The results will also be important for the interpretation of the spectra of suspected and newly identified binary and multiple systems.


Significant Spacecraft Anomalies: (The following are preliminary reports of potential non-nominal performance that will be investigated.)


11539 – ACS suspended during High Voltage turn on and SBC transition

ACS suspended with status buffer messages ACS 715, parameter 166, and ACS 707, parameter 0, both occurring at time 61165 corresponding to a time of 17:40:16.625. Solar Blind Camera MAMA High Voltage turn on was in progress at that time.

11540 – NSSC-1 safed.

Upon acquisition of signal at 290/21:17:11 the NSSC-1 had safed. Two 486 ESB messages “D24” (SICDH_TOGGLE_TEST_FAIL) and “D01″(SAFE_HLD_MACROS_ACTIVE_INFO) were received at 290/21:14:29. SCDHSAFA (Safing Macro Active), SDSTOGF (Toggle Failure Counter), SSiCDHT (SIC&DH Toggle OK Flag), and SSIPTBE (PIT Toggle Test) flags were set.


18329-0 – NSSC-1 Pointer for Science SMS Intercept @ 290/1406z
18332-3 – Dump ACS Memory @ 290/1939z
18333-1 – NSSC-1 HW Memory Dump @ 290/2213z


NS-20 – WFPC-II Decontamination Limits Management @290/1746z


FGS GSacq               06                   06
FGS REacq               01                   01
OBAD with Maneuver      14                   14


Flash Report :

HST Status as of Thursday evening 291/00:00: At 290/17:40:16 UTC, ACS suspended due to the Solar Blind Camera MAMA low voltage power supply apparently not coming up properly. At 290/21:14 the NSSC-1 halted. The halting resulted in the HST486 SIC&DH PIT toggle test failing twenty seconds later safing the payload with the execution of SMAC20.

SpaceRef staff editor.